40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 10 12:42:27 2018, gautam, Update, SUS, ETMX DC alignment ETMXdrift.png
    Reply  Wed Jan 10 18:53:31 2018, gautam, Update, SUS, ETMX DC alignment ETMXdrift.png
       Reply  Wed Jan 10 22:19:44 2018, rana, Update, SUS, ETMX DC alignment 
Message ID: 13523     Entry time: Wed Jan 10 12:42:27 2018     Reply to this: 13527
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMX DC alignment 

I've been observing this for a few days: ETMX's DC alignment seems to drift by so much that the previously well aligned X arm cavity is now totally misaligned.

The wall StripTool trace shows that both the X and Y arms were locked with arm transmissions around 1 till c1psl conked out - so in the attached plot, around 1400 UTC, the arm cavity was well aligned. So the sudden jump in the OSEM sensor signals is the time at which LSC control to the ETM was triggered OFF. But as seen in the attached plot, after the lockloss, the Oplev signals seem to show that the mirror alignment drifted by >50urad. This level of drift isn't consistent with the OSEM sensor signals - of course, the Oplev calibration could be off, but the tension in values is almost an order of magnitude. The misalignment seems real - the other Oplev spots have stuck around near the (0,0) points where I recentered them last night, only ETMX seems to have undergone misalignment.

Need to think about what's happening here. Note that this kind of "drift" behaviour seems to be distinct from the infamous ETMX "glitching" problem that was supposed to have been fixed in the 2016 vent.

 

Attachment 1: ETMXdrift.png  66 kB  | Hide | Hide all
ETMXdrift.png
ELOG V3.1.3-